# Exploring Decision-based Black-box Attacks on Face Forgery Detection

### v1

@@ Abstract
Face forgery generation technologies generate vivid faces, which have raised public concerns about security and privacy. Many intelligent systems, such as electronic payment and identity verification, rely on face forgery detection. Although face forgery detection has successfully distinguished fake faces, recent studies have demonstrated that face forgery detectors are very vulnerable to adversarial examples. Meanwhile, existing attacks rely on network architectures or training datasets instead of the predicted labels, which leads to a gap in attacking deployed applications. To narrow this gap, we first explore the decision-based attacks on face forgery detection. However, applying existing decision-based attacks directly suffers from perturbation initialization failure and low image quality. First, we propose cross-task perturbation to handle initialization failures by utilizing the high correlation of face features on different tasks. Then, inspired by using frequency cues by face forgery detection, we propose the frequency decision-based attack. We add perturbations in the frequency domain and then constrain the visual quality in the spatial domain. Finally, extensive experiments demonstrate that our method achieves state-of-the-art attack performance on FaceForensics++, CelebDF, and industrial APIs, with high query efficiency and guaranteed image quality. Further, the fake faces by our method can pass face forgery detection and face recognition, which exposes the security problems of face forgery detectors.
Index Terms:
Adversarial examples, Face forgery detection, Black-box attacks, Face recognition
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
With the success of deep neural networks [1, 2, 3, 4, 5] and the development of generative models [6, 7, 8, 9], face forgery generation has made excellent progress. We can synthesize highly realistic fake faces by these techniques, such as Face2Face [10], FaceSwap [11], DeepFakes [12] and NeuralTextures [13]. Further, methods such as StyleGAN [14] edit one’s attributes (hair, glasses, age, etc.) to generate various faces, and these faces are widely used in scenes such as the film industry and virtual uploaders. In contrast, these fake faces also lead to the spread of false news or damage to reputation, raising public concerns about security and privacy [15]. Therefore, researchers have recently focused on how to design effective face forgery detection methods to determine whether a face has been modified.
@@ I Introduction
Face forgery detection is usually defined as a binary classification problem, distinguishing between real and fake faces. There is a series of face forgery detectors [1, 2, 3, 4, 16, 17] proposed that utilize deep neural networks to automatically learn feature differences between real and fake faces and achieve remarkable success on multiple datasets [18, 19]. Despite their success, recent work [20] shows that face forgery detectors are very vulnerable to adversarial examples [21, 22, 23], which exposes the current shortcut of security in face forgery detection. After adding a carefully-designed adversarial perturbation, a face originally classified as the fake is recognized as real by face forgery detectors. Existing methods [20, 24, 25, 26, 27, 28] have explored the adversarial robustness of face forgery detection, but these methods require access to the detector’s network architecture or training datasets instead of the predicted labels. However, when attacking industrial APIs, we cannot access the details of the models or datasets, but only the decision (real or fake) returned by the service. Therefore, the lack of decision-based attacks leads to a gap in the robustness analysis of face forgery detection.
@@
When exploring decision-based attacks on face forgery detection, it is easiest to apply existing decision-based attacks directly [29, 30, 31, 32, 33, 34]. However, directly applying existing decision-based attacks to face forgery detection may encounter some issues: (a) perturbation initialization failure. Existing decision-based attacks usually adopt random noise as attack initialization. However, since adding noise itself is a kind of manipulation, faces with random noise are usually recognized as fake faces by face forgery detection (as shown in Fig. 1). (b) low image quality. For example, RayS [35] introduces perceptible regular rectangular perturbation. Using real faces as perturbation initialization [33] leads to artifacts and may affect the application of faces, e.g. face recognition, as shown in Fig. 1. Therefore, we need an efficient decision-based attack designed for face forgery detection to identify potential vulnerabilities on face forgery detection and provide insights on possible improvements.
@@
To address these issues, we propose an efficient decision-based adversarial attack to explore the adversarial robustness of face forgery detection. First, we propose a cross-task perturbed initialization method to handle initialization failures of decision-based attacks on face forgery detection. Specifically, although it is difficult for us to obtain face forgery detectors because we do not know the face and its modification method, it is relatively easy to obtain many open-source face recognition models. With the high correlation of face features on two types of tasks, we generate cross-task adversarial perturbations in the face recognition model as initial perturbations to attack face forgery detection. Second, since the features of face forgery detection are highly discriminative in the frequency domain [36, 16, 17], we propose a frequency decision-based attack. After obtaining cross-task adversarial perturbations, we add perturbations to faces in the frequency domain and then constrain the visual quality in the spatial domain. Finally, our method achieves state-of-the-art attack performance on FaceForensics++ [18] and CelebDF [19], with high query efficiency and guaranteed image quality. We also successfully attack the real-world industrial API and showcase the effectiveness of our method in practice. Furthermore, fake faces by our method can simultaneously pass face forgery detection and face recognition, which exposes the current adversarial vulnerability of face-related systems.
@@
  • We first explore decision-based attacks on face forgery detection and verify their adversarial vulnerability.
@@
  • Inspired by the frequency clues of face forgery detection, we propose a frequency decision-based attack that adds perturbations in the frequency domain and then constrains the visual quality in the spatial domain.
  • Experiments show that our method achieves state-of-the-art attack performance on FaceForensics++, CelebDF, and industrial APIs, which exposes the vulnerability of face forgery detection under decision-based attacks.
@@
The remainder of the paper is organized as follows. Section II briefly introduces face forgery detection and decision-based black-box attacks, and then discusses adversarial attacks on face forgery detection. Section III first gives the formal definition of decision-based attacks on face forgery detection. Then, we introduce Cross-task Perturbation Initialization and propose the Frequency Decision-based Attack. Section IV presents experiments to illustrate the effectiveness of the proposed method, including attacks on state-of-the-art detectors, ablation studies, attacks on industrial APIs, etc. We further discuss the possibility of using face-related tasks to assist in attacking face forgery detection in Section V and conclude this paper in Section VI.
II Related Work
II-A Face Forgery Detection
[…]
II-B Decision-based Black-box Attacks
@@
In contrast to white-box attacks that require access to model details such as architecture and gradients, recent work has focused on black-box attacks [29], including transfer-based, score-based, and decision-based attacks. Transfer-based attacks need to know the distribution of the data to construct the surrogate model and its performance is limited by the surrogate model. The difference between score-based and decision-based attacks is that the output returned by the threat model is a confidence or a label. It is obvious that the decision-based setting is the most difficult, but it is also the most practical since in reality, the adversary only knows minimal information about the model.
@@
Brendel et al. [29] propose the first decision-based attack to randomly walk on the decision boundary starting from an adversarial point while keeping adversarial, called the Boundary Attack. Some works [30, 31, 34, 32] improve Boundary Attack from the perspective of gradient optimization. Next, SignFlip [33] introduces noise projection and random sign flip to improve decision-based attack by large margins. RayS [35] reduces the number of queries by reformulating the continuous problem of finding the closest decision boundary into a discrete problem, which does not require zeroth-order gradient estimation. In this paper, we explore decision-based attacks on face forgery detection under $l_{\infty}$ norm, i.e. adding perturbation within the perturbation range $\epsilon$, which is more practical in real-world applications.
II-C Adversarial Attacks on Face Forgery Detection
@@
Face forgery detection is a security-sensitive task and its adversarial vulnerability has received much attention [20, 24, 25, 26, 27]. Carlini et al. [20] and Neekhara et al. [25] shows the adversarial robustness of face forgery detectors under various white-box attacks and transfer-based attacks. Hussain et al. [37] evaluate the detectors by the score-based attack with natural evolutionary strategies [38]. Li et al. [26] manipulate the noise and latent vectors of StyleGAN [14] to fool face forgery detectors. KRA [24] attacks the key regions of forged faces obtained from the semantic region selection and boosts the adversarial transferability. Then, Jia et al. [27] add adversarial perturbations in the frequency domain and generate more imperceptible adversarial examples with a high transferability. In this paper, considering academic and industrial applications, we focus on attacking frame-level detectors because recent attacks on face forgery detection have mainly focused on attacking frame-level detectors [20, 24, 25, 26, 27], and industrial APIs tend to use images as input because of the low computation and communication costs. In conclusion, we evaluate the adversarial robustness of face forgery detection under decision-based attacks for the first time and hope to provide insight into possible future improvements.
III Methodology
@@
III-A Preliminaries
We define a neural network-based face forgery detector $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ as the threat model. Because the face forgery detection is a binary classification task, its label $y$ is real or fake, i.e. $y\in\mathbb{R}^{2}=\{0,1\}$. Given a face $x\in[0,1]^{d}$, we define $f(x)_{i}$ as the probability of class $i$ and $c(x)=\mathrm{argmax}_{i\in\{0,1\}}\ f(x)_{i}$ refers to the predicted label. In this paper, we only consider faces that are classified as fake and treat the situation where fake faces are classified as real faces as successful attacks. Therefore, the adversary aims to find an adversarial perturbation $\delta\in\mathbb{R}^{d}$ such that $c(x+\delta)=0$ (classified as real). Here, we define $||\delta||_{\infty}\leq\epsilon$ and $\epsilon$ is the allowed maximum perturbation.
@@ III-A Preliminaries
The adversary keeps querying the model and optimizing adversarial perturbations $\delta$ until $||\delta||_{\infty}\leq\epsilon$.
Figure 2: Empirical analysis of cosine similarity (%) on intermediate features between forgery detectors and face recognition models under clean examples and adversarial examples (FGSM [22], PGD [39], and MIM [40]). We find that face recognition models and face forgery detectors have a high correlation between the intermediate features of the same face. It means that the perturbation of attacking face recognition can also perturb face forgery detection to a certain extent.
Figure 3: (a) The overview of cross-task perturbation initialization. With the high correlation of intermediate features between face-related tasks, we iteratively update the real face, keeping it adversarial while improving efficiency for subsequent attacks. (b) The pipeline of the frequency decision-based attack. We first obtain the initial perturbation by perturbation initialization. Then, we iterate in frequency noise projection and random perturbation flip until $c(x_{adv})=0$ and $||\delta||_{\infty}\leq\epsilon$.
III-B Cross-task Perturbation Initialization
The easiest way to explore decision-based attacks in face forgery detection is to apply existing decision-based attacks directly [29, 30, 31, 32, 33, 34]. In the case study, we choose the state-of-the-art HSJA [32], SignFlip [33], and RayS [35]. We apply these attacks directly to face forgery detection and find that they cause the attack to fail. Specifically, they all have a situation where the initialization fails, and the attack success rate is close to 0%. It shows that the random initialization is unsuccessful and cannot guarantee adversarial. Face forgery detection usually recognizes faces with random noises as fake faces because adding random noises is a common manipulation. Furthermore, using real faces as initialization can guarantee adversarial, but it may lead to artifacts, so image quality is not guaranteed. Hence, we hope a perturbation initialization can handle this issue while ensuring query efficiency.
Algorithm 1 Cross-task Perturbation Initialization

### v2

@@ Abstract
Face forgery generation technologies generate vivid faces, which have raised public concerns about security and privacy. Many intelligent systems, such as electronic payment and identity verification, rely on face forgery detection. Although face forgery detection has successfully distinguished fake faces, recent studies have demonstrated that face forgery detectors are very vulnerable to adversarial examples. Meanwhile, existing attacks rely on network architectures or training datasets instead of the predicted labels, which leads to a gap in attacking deployed applications. To narrow this gap, we first explore the decision-based attacks on face forgery detection. We identify challenges in directly applying existing decision-based attacks, such as perturbation initialization failure and reduced image quality. To overcome these issues, we propose cross-task perturbation to handle initialization failures by utilizing the high correlation of face features on different tasks. Additionally, inspired by the use of frequency cues in face forgery detection, we introduce the frequency decision-based attack. This attack involves adding perturbations in the frequency domain while constraining visual quality in the spatial domain. Finally, extensive experiments demonstrate that our method achieves state-of-the-art attack performance on FaceForensics++, CelebDF, and industrial APIs, with high query efficiency and guaranteed image quality. Further, the fake faces by our method can pass face forgery detection and face recognition, which exposes the security problems of face forgery detectors.
keywords:
[…]
1. Introduction
With the success of deep neural networks [resnet, xception, EfficientNet, vit, chen2022towards] and the development of generative models [DBLP:conf/nips/GoodfellowPMXWOCB14, ddim, beatgan, ldm], face forgery generation has made excellent progress. We can synthesize highly realistic fake faces by these techniques, such as Face2Face [Face2Face], FaceSwap [Faceswap], DeepFakes [Deepfake] and NeuralTextures [neuraltextures]. Further, methods such as StyleGAN [stylegan] edit one’s attributes (hair, glasses, age, etc.) to generate various faces, and these faces are widely used in scenes such as the film industry and virtual uploaders. In contrast, these fake faces also lead to the spread of false news or damage to reputation, raising public concerns about security and privacy [sun2023contrastive, fas1, fas2, fas3, fas4, fas5]. Therefore, researchers have recently focused on how to design effective face forgery detection methods to determine whether a face has been modified.
@@ 1. Introduction
Face forgery detection is usually defined as a binary classification problem, distinguishing between real and fake faces. There is a series of face forgery detectors [resnet, xception, EfficientNet, vit, f3net, M2TR] proposed that utilize deep neural networks to automatically learn feature differences between real and fake faces and achieve remarkable success on multiple datasets [ffpp, CelebDF]. Despite their success, recent work [EvadingDF] shows that face forgery detectors are very vulnerable to adversarial examples [chen2022shape, fgsm, devopatch], which exposes the current shortcut of security in face forgery detection. After adding a carefully-designed adversarial perturbation, a face originally classified as the fake is recognized as real by face forgery detectors. Existing methods [EvadingDF, KeyRegionAttack, AdvDfPractical, FaceManifold, hybridfre, huang2022cmua] have explored the adversarial robustness of face forgery detection, but these methods require access to the detector’s network architecture or training datasets. When targeting industrial APIs in the context of face forgery detection, we do not have access to the underlying models or datasets. Instead, we can only observe the binary decision (real or fake) provided by the service. Consequently, the absence of decision-based attacks results in a significant gap in the robustness evaluation of face forgery detection systems.
@@
When exploring decision-based attacks on face forgery detection, it is easiest to apply existing decision-based attacks directly [boundary, OptAttack, signopt, hsja, SignFlip, TA]. However, directly applying existing decision-based attacks to face forgery detection may encounter some issues: (a) perturbation initialization failure. Existing decision-based attacks usually adopt random noise as attack initialization. However, since adding large noise itself is a kind of manipulation, faces with large random noise are usually recognized as fake faces by face forgery detection (as shown in Figure 1). (b) low image quality. For example, RayS [RayS] introduces perceptible regular rectangular perturbation. Using real faces for perturbation initialization [SignFlip] can introduce artifacts and potentially impact face-related applications, such as face recognition. Even if such faces succeed in bypassing forgery detection models, they fail to achieve the ultimate goal of deceiving face recognition systems, as illustrated in Figure 1. Therefore, we need an efficient decision-based attack designed for face forgery detection to identify potential vulnerabilities on face forgery detection and provide insights on possible improvements.
@@
To address these issues, we propose an efficient decision-based adversarial attack to explore the adversarial robustness of face forgery detection. First, we propose a cross-task perturbation initialization method to handle initialization failures of decision-based attacks on face forgery detection. With the high correlation of face features on face forgery detection and face recognition, we generate cross-task adversarial perturbations in the face recognition model as initial perturbations to attack face forgery detection. Second, since the features of face forgery detection are highly discriminative in the frequency domain [fre_analysis, f3net, M2TR], we propose a frequency decision-based attack. After obtaining cross-task adversarial perturbations, we add perturbations to faces in the frequency domain and then constrain the visual quality in the spatial domain. Finally, our method achieves state-of-the-art attack performance on FaceForensics++ [ffpp] and CelebDF [CelebDF], with high query efficiency and guaranteed image quality. We also successfully attack the real-world industrial API and showcase the effectiveness of our method in practice. Furthermore, fake faces by our method can simultaneously pass face forgery detection and face recognition, which exposes the current adversarial vulnerability of face-related systems.
@@
  • We first explore decision-based attacks on face forgery detection and verify their adversarial vulnerability in academic and industrial scenarios. Fake faces by our method can pass face forgery detection and face recognition, which further exposes the security problems of face forgery detectors.
@@
  • We propose a frequency decision-based attack based on frequency clues, which for the first time verifies the effectiveness of frequency domain-based query attacks on face forgery detection.
  • Experiments show that our method achieves state-of-the-art attack performance on FaceForensics++ and CelebDF, and first show that frequency-based detectors are just as vulnerable as spatial-based ones.
@@
The remainder of the paper is organized as follows. Section 2 briefly introduces face forgery detection and decision-based black-box attacks and then discusses adversarial attacks on face forgery detection. Section 3 first gives the formal definition of decision-based attacks on face forgery detection. Then, we introduce Cross-task Perturbation Initialization and propose the Frequency Decision-based Attack. Section 4 presents experiments to illustrate the effectiveness of the proposed method, including attacks on state-of-the-art detectors, ablation studies, attacks on industrial APIs, etc. We further discuss the possibility of using face-related tasks to assist in attacking face forgery detection in Section 5 and conclude this paper in Section 6.
2. Related Work
2.1. Face Forgery Detection
Since large-scale public datasets are freely available, combined with rapid progress in generative models, a very realistic fake face has been generated with its corresponding implications for society. Therefore, there have been many efforts to distinguish real and fake faces as much as possible in face forgery detection. Early work [ffpp] directly uses convolutional neural networks (CNNs) to extract discriminative features for face forgery detection. However, these spatial features extracted by CNNs focus more on class-level distinctions rather than subtle differences between real and fake faces. Furthermore, recent work [fre_analysis, f3net, M2TR] mines forgery patterns and proposes many frequency-based detection methods, noting that real and fake faces are diverse in the frequency domain. As a two-branch network, F3Net [f3net] finds forgery patterns using frequency clues and extracts discrepancies between real and fake images by looking at frequency statistics. Considering the potential of the recent Vision Transformer [vit], M2TR [M2TR] utilizes a multi-scale transformer for capturing local inconsistencies at different scales and additionally introduces frequency modality to improve performance with multiple image compression algorithms. In addition, recent works have utilized self-augmentation [sbi], attention [fed2023, nguyen2024laa], multimodality [modalffd], and mixture-of-experts [moeffd] to further obtain more generalized detectors. To better evaluate the adversarial robustness of face forgery detection, we choose more classic and more generalizable spatial-based [resnet, xception, EfficientNet, vit] and frequency-based detectors [f3net, M2TR] for experiments.
2.2. Decision-based Black-box Attacks
@@
In contrast to white-box attacks that require access to model details such as architecture and gradients, recent work has focused on black-box attacks [boundary, chen2023aca, Jiang_2023_ICCV, 10.1145/3581783.3611828, WANG2024124757, 10.1145/3581783.3611827], including transfer-based, score-based, and decision-based attacks. Transfer-based attacks need to know the distribution of the data to construct the surrogate model and its performance is limited by the surrogate model. The difference between score-based and decision-based attacks is that the output returned by the threat model is a confidence or a label. It is obvious that the decision-based setting is the most difficult, but it is also the most practical since in reality, the adversary only knows minimal information about the model.
@@
Brendel et al. [boundary] propose the first decision-based attack to randomly walk on the decision boundary starting from an adversarial point while keeping adversarial, called the Boundary Attack. Some works [OptAttack, signopt, TA, hsja] improve Boundary Attack from the perspective of gradient optimization. Next, SignFlip [SignFlip] introduces noise projection and random sign flip to improve decision-based attack by large margins. RayS [RayS] reduces the number of queries by reformulating the continuous problem of finding the closest decision boundary into a discrete problem, which does not require zeroth-order gradient estimation. In this paper, we explore decision-based attacks on face forgery detection under $l_{\infty}$ norm, i.e. adding perturbation within the perturbation range $\epsilon$, which is more practical in real-world applications.
2.3. Adversarial Attacks on Face Forgery Detection
@@
Face forgery detection is a security-sensitive task and its adversarial vulnerability has received much attention [EvadingDF, KeyRegionAttack, AdvDfPractical, FaceManifold, hybridfre]. Carlini et al. [EvadingDF] and Neekhara et al. [AdvDfPractical] shows the adversarial robustness of face forgery detectors under various white-box attacks and transfer-based attacks. Hussain et al. [advdeepfake] evaluate the detectors by the score-based attack with natural evolutionary strategies [nes]. Li et al. [FaceManifold] manipulate the noise and latent vectors of StyleGAN [stylegan] to fool face forgery detectors. KRA [KeyRegionAttack] attacks the key regions of forged faces obtained from the semantic region selection and boosts the adversarial transferability. Then, Jia et al. [hybridfre] add adversarial perturbations in the frequency domain and generate more imperceptible adversarial examples with a high transferability. However, existing work is unable to evaluate adversarial robustness in a decision-based setting. In this paper, considering academic and industrial applications, we focus on attacking frame-level detectors because recent attacks on face forgery detection have mainly focused on attacking frame-level detectors [EvadingDF, KeyRegionAttack, AdvDfPractical, FaceManifold, hybridfre], and industrial APIs tend to use images as input because of the low computation and communication costs. In conclusion, we evaluate the adversarial robustness of face forgery detection under decision-based attacks for the first time and hope to provide insight into possible future improvements.
3. Methodology
@@
3.1. Preliminaries
We define a neural network-based face forgery detector $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ as the threat model. Because the face forgery detection is a binary classification task, its label $y$ is real or fake, i.e. $y\in\mathbb{R}^{2}=\{0,1\}$. Given a face $x\in[0,1]^{d}$, we define $f(x)_{i}$ as the probability of class $i$ and $c(x)=\mathrm{argmax}_{i\in\{0,1\}}\ f(x)_{i}$ refers to the predicted label. In this paper, we only consider faces that are classified as fake and treat the situation where fake faces are classified as real faces as successful attacks. Therefore, the adversary aims to find an adversarial perturbation $\delta\in\mathbb{R}^{d}$ such that $c(x+\delta)=0$ (classified as real). Here, we define $\delta||_{\infty}\leq\epsilon$ and $\epsilon$ is the allowed maximum perturbation.
@@ 3.1. Preliminaries
The adversary keeps querying the model and optimizing adversarial perturbations $\delta$ until $\delta||_{\infty}\leq\epsilon$.
Figure 2: Empirical analysis of cosine similarity (%) on intermediate features between forgery detectors and face recognition models under clean examples and adversarial examples (FGSM [fgsm], PGD [pgd], and MIM [mim]). Here, the face recognition models are FaceNet [FaceNet], CosFace [CosFace], and ArcFace [ArcFace]. The face forgery detection models are ResNet50 [resnet], Xception [xception], EfficientNet-b4 (Eb4) [EfficientNet]. We find that face recognition models and face forgery detectors have a high correlation between the intermediate features of the same face. It means that the perturbation of attacking face recognition can also perturb face forgery detection to a certain extent.
Figure 3: (a) The overview of cross-task perturbation initialization. With the high correlation of intermediate features between face-related tasks, we iteratively update the real face, keeping it adversarial while improving efficiency for subsequent attacks. (b) The pipeline of the frequency decision-based attack. We first obtain the initial perturbation by perturbation initialization. Then, we iterate in frequency noise projection and random perturbation flip until $c(x_{adv})=0$ and $\delta||_{\infty}\leq\epsilon$.
3.2. Cross-task Perturbation Initialization
In this section, we introduce the cross-task perturbation initialization (CPI) to handle the attack initialization failure. Actually, the easiest way to explore decision-based attacks in face forgery detection is to apply existing decision-based attacks directly [boundary, OptAttack, signopt, hsja, SignFlip, TA]. In the case study, we choose the state-of-the-art HSJA [hsja], SignFlip [SignFlip], and RayS [RayS]. We apply these attacks directly to face forgery detection and find that they cause the attack to fail. Specifically, they all have a situation where the initialization fails, and the attack success rate is close to 0%. It shows that the random initialization is unsuccessful and cannot guarantee adversarial, as shown in the random initialization in Figure 1. Face forgery detection usually recognizes faces with large random noises as fake faces because adding large random noises is a common manipulation. In addition, due to the large random noise, such samples usually fail face recognition. Furthermore, using real faces as initialization can guarantee adversarial, but it may lead to artifacts, so image quality is not guaranteed. Hence, we hope a perturbation initialization can handle this issue while ensuring query efficiency.
Naturally, exploiting the adversarial transferability of face forgery detectors is a possible strategy. However, in practice, we not only do not have access to the training data of the threat model but also do not know the face forgery generation method it adopts, which greatly limits the possibility of using transferable adversarial perturbations. Since face recognition is a very common face-related task and its open-source model is relatively easy to obtain, it may help the attack achieve perturbation initialization. Furthermore, intermediate features may have high correlations between different face-related tasks. From this perspective, using face recognition models to transfer adversarial perturbations is a feasible initialization strategy.
Algorithm 1 Cross-task Perturbation Initialization (CPI)